I know that the table represents a linear function, but I'm having difficulty coming up the linear equation that models the data.
x: 0 5 10 15
f(x): -6 9 24 39
My answers so far have been: -3x+5, 3x+5, 5x+3, and 5x+ 15.

Answer by Theo(13342)

looks like the slope is 3.

change in y divided by change in x = 15/5 = 3

y-intercept is -6 (value of y when value of x is equal to 0.

equation should be y = 3x - 6

when x = 0, y = -6
when x = 5, y = 15 - 6 = 9
when x = 10, y = 30 - 6 = 24
when x = 15, y = 45 - 6 = 39

that's your solution.

slope intercept form of the equation of a straight line is y = mx + b

m is the slope
b is the y-intercept

slope is change in y / change in x.

your data fits the equation of y = 3x - 6.

here's a graph of the equation.

$$$